OpenStudy (anonymous):

7x+4=9y/4, solve for y

OpenStudy (anonymous):

7x+4=9y/4 4(7x+4)=9y (28x+16)/9=y

OpenStudy (anonymous):

y = (28x+16)/9

OpenStudy (anonymous):

y=(28x+16)/9